# The 4th Deep Learning for Code Workshop

### v2

@@ Appendix C Encoder Details
Table 4: Hyperparameters of the embedder training
The hyperparameters used for training the embedder model are listed in Table˜4. We have used microsoft/codebert-base as the base model and trained our embedder for 22000 steps with a batch size 16, since we use many negative samples in the loss. We applied a dropout of 0.1 on the last layer of the model; the embedding dimension is 768, and the maximum sequence length is 128. We use AdamW optimizer with a linear warmup schedule for 10% of the training steps.
C.1 Visualizing RocqStar vs. Baseline Premise Selection
[…]
Proof using.
[…]
  by rewrite H2.
Qed.
[…]
Proof using.
[…]
  lia.
Qed.
Figure 3: Theorems with dissimilar statements and similar proofs
$\operatorname{Jaccard\_distance}(t1,t2)$ $=1-\frac{\left|\{{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{transitive}},{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{ext\_sb}}\}\cap\{{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{irreflexive}},{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{ext\_sb}}\}\right|}{\left|\{{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{transitive}},{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{ext\_sb}}\}\cup\{{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{irreflexive}},{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{ext\_sb}}\}\right|}$
$=1-\frac{1}{3}=0.67$
[…]
$\operatorname{proof\_sim}(t1,t2)$ $=0.32$
$\operatorname{embedder\_pred}(t1,t2)$ $=0.28$
Appendix D Agent toolset
@@
Appendix E Plan refinement via Multi-Agent Debate
We will illustrate the difference between the two plans; one was generated using a single call to a model, and the other was generated using multi-agent debate. The theorem we are trying to prove is the sub_IPROP from the src/imm/SubExecution.v file, that states:
@@ Appendix E Plan refinement via Multi-Agent Debate
The following plan was produced on the planning stage via a single request to the model, without multi-agent debate:
@@
rewrite sub_W.
rewrite sub_is_ta_propagate_to_G.
@@
3. Reasoning Steps: Work with explicit set inclusion based on subseteq definition Use any discovered lemmas about G/G’ relationship Preserve flexibility to adapt based on actual definitions
@@
Appendix F Experiments compute resources
Agent evaluation  In the case of the agent evaluation, we ran the experiment only once and did not provide the confidence intervals due to financial limitations. We ran our agent on the IMM-300 dataset, and afterward, we compared five versions of the agent on the IMM-50 dataset in our §˜4.3. That results in 550 attempts to prove different theorems. We estimate a single attempt at 1.3 US dollars. Therefore, we estimate the cost of the evaluation of the agent to be 715 US dollars.

### v3

@@ Appendix B Encoder Details
Table 4. Hyperparameters of the embedder training
The hyperparameters used for training the embedder model are listed in Table 4. We have used microsoft/codebert-base as the base model and trained our embedder for 22000 steps with a batch size 16, since we use many negative samples in the loss. We applied a dropout of 0.1 on the last layer of the model; the embedding dimension is 768, and the maximum sequence length is 128. We use AdamW optimizer with a linear warmup schedule for 10% of the training steps.
Appendix C Agent toolset
@@
Appendix D IMM Evaluation Dataset
The collected IMM-300 dataset from the CoqPilot coqpilot work includes only theorems with proofs of length no more than 20. For that reason the bucket with the most difficult theorems is labeled 9—20 tactics. This decision has been made, reflecting CoqPilot’s original focus on subgoals and shorter lemmas. Theorems of length no more than 20 tactics account for 83% of all proofs in the IMM project. As we take the same dataset, it possesses the same limitations. Therefore, we have not evaluated our solution on theorems, for which the reference proof contains more than 20 tactics. However, such theorems are quite rare.
The exact list of theorems used in each group could be found in the repository of the CoqPilot project: https://github.com/JetBrains-Research/coqpilot/blob/main/etc/docs/benchmark/.
A common problem with testing pipelines that include general-purpose LLM providers, such as OpenAI, is data contamination. We are aware, that the model could have possibly seen the human-written proofs, as the IMM project has been publicly available since a while. However, firstly, the model is provided with neither the theorem name for which it is generating the proof, nor the proof goal exactly as it appeared in the original file. As we treat them as proof states, rather than theorems, an LLM receives it in an equivalent, but slightly modified way. Secondly, as many of our experiments have shown, various quality of premise selection drasticly changes the behavior of the model. That hints that the model is not able to memorize all theorems and proofs. Lastly, data contamination issue was one of the things we had in mind, while developing BigRocq. One could pass a Rocq project into BigRocq as input, and for each theorem retrieve the sub-state, that is achieved after $k$ steps. On an example of $k=2$, the following theorem:
Lemma eq_trans (A : Type) : forall (x y z : A), x = y -> y = z -> x = z.
Proof.
intros x y z Hxy Hyz.
rewrite Hxy. (* State: (A : Type) (x y z: A) (Hxy: x = y) (Hyz: y = z) : y = z *)
rewrite Hyz.
reflexivity.
Qed.
Could be automatically tranformed into the following one:
Proof.
rewrite Hyz.
reflexivity.
Qed.
D.1. Visualizing RocqStar vs. Baseline Premise Selection
[…]
Proof using.
[…]
  by rewrite H2.
Qed.
[…]
Proof using.
[…]
  lia.
Qed.
Figure 3. Theorems with dissimilar statements and similar proofs
$\operatorname{Jaccard\_dist}(t1,t2)$ $=1-\frac{\left|\{{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{transitive}},{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{ext\_sb}}\}\cap\{{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{irreflexive}},{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{ext\_sb}}\}\right|}{\left|\{{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{transitive}},{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{ext\_sb}}\}\cup\{{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{irreflexive}},{\color[rgb]{0.0703125,0.3671875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.3671875,0.1328125}\texttt{ext\_sb}}\}\right|}$
$=1-\frac{1}{3}=0.67$
[…]
$\operatorname{proof\_sim}(t1,t2)$ $=0.32$
$\operatorname{embedder\_pred}(t1,t2)$ $=0.28$
Appendix E Experiments compute resources
Agent evaluation  In the case of the agent evaluation, we ran the experiment only once and did not provide the confidence intervals due to financial limitations. We ran our agent on the IMM-300 dataset, and afterward, we compared five versions of the agent on the IMM-50 dataset in our § 4.3. That results in 550 attempts to prove different theorems. We estimate a single attempt at 1.3 US dollars. Therefore, we estimate the cost of the evaluation of the agent to be 715 US dollars.
Appendix F Plan refinement via Multi-Agent Debate
We compare two proof plans: one generated by a single model call and one produced via multi-agent debate. The target theorem is sub_IPROP from src/imm/SubExecution.v, which states:
@@ Appendix F Plan refinement via Multi-Agent Debate
The following plan was produced by a single planning call:
@@
rewrite sub_W. rewrite sub_is_ta_propagate_to_G.
@@
(3) Reasoning Steps: Work with explicit set inclusion based on subseteq definition Use any discovered lemmas about G/G’ relationship Preserve flexibility to adapt based on actual definitions
@@
Appendix G Agent Execution Patterns
Average successful run
Average failing run
Figure 4. Distributions of successful and failing runs for the no_sb_cr_to_init theorem
We analyze the agent’s behavior to identify patterns that distinguish successful and failed proof attempts. Specifically, we selected the theorem no_sb_cr_to_init from group B of the IMM-50 dataset, which the agent sometimes proved and sometimes failed to prove. We collected traces from ten successful and seventeen failing executions and plotted the distributions of tool calls averaged over all runs in each scenario (Fig. 4).
Although the experiment was relatively small, it already provides several observations about the agent’s behavior. We have not yet identified clear action points that could directly improve the agent, but we plan to conduct a larger-scale study and dedicate more effort to analyzing execution trajectories in the future. Notably, only a subset of available tools is actively used during proof attempts, especially in successful runs, indicating that the agent has not yet learned to leverage the full range of available functionality, although this behavior partly reflects the specifics of the analyzed theorem. In failing runs, we observe a more uniform distribution of tool calls (particularly aboutTerm, printTerm, and searchPattern), without a pronounced early activity peak. This behavior suggests that the agent tends to wander through the search space without a clear strategy. Combined with the findings from § 4.3, which highlight the importance of effective planning, this supports the view that a well-defined plan, one that explicitly incorporates context exploration and the retrieval of relevant information before trying to build the proof, is essential for stable performance. We consider introducing a dedicated, more manual phase for contextual exploration in future iterations of the agent.